# Diversity-grounded Channel Prototypical Learning for Out-of-Distribution Intent Detection

### v1

@@ Abstract
Bo Liu1, Liming Zhan1, Yujie Feng1, Zexin Lu1, Chengqiang Xie1, Lei Xue2, Xiao-Ming Wu1thanks:   Corresponding author., Albert Y.S. Lam3 1Department of Computing, The Hong Kong Polytechnic University, Hong Kong S.A.R. 2School of Cyber Science and Technology, Sun Yat-Sen University, Shenzhen, China 3Fano Labs, Hong Kong S.A.R. bokelvin.liu@connect.polyu.edu.hk, csxmwu@comp.polyu.edu.hk
@@ 1 Introduction
Recent advancements in large language models (LLMs) Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023) have significantly improved the detection of semantically distinct out-of-distribution (far-OOD) intents Liu et al. (2023). However, identifying semantically similar (near-OOD) intents continues to pose significant challenges Fort et al. (2021); Liu et al. (2023). Given its critical importance in practical applications, our study concentrates on near-OOD detection, specifically in a few-shot scenario Jiang et al. (2024); Zhang et al. (2022b). Particularly, in the few-shot learning context, the training model can receive only a few in-distribution (ID) examples per class. Obtaining discriminative information from limited ID examples for OOD detection is inherently challenging Zhang et al. (2020a). Previous strategies have involved using extra unlabeled data Wang et al. (2023) or generating pseudo-ID examples Zhan et al. (2022) to increase training samples. Instead, we focus on leveraging the inner knowledge of LLMs as a supplement without dataset expansion.

### v2

@@ Abstract
Bo Liu1, Liming Zhan1, Yujie Feng1, Zexin Lu1, Chengqiang Xie1, Lei Xue2, Albert Y.S. Lam3, Xiao-Ming Wu1thanks:   Corresponding author. 1Department of Computing, The Hong Kong Polytechnic University, Hong Kong S.A.R. 2School of Cyber Science and Technology, Sun Yat-Sen University, Shenzhen, China 3Fano Labs, Hong Kong S.A.R. bokelvin.liu@connect.polyu.edu.hk, csxmwu@comp.polyu.edu.hk
@@ 1 Introduction
Recent advancements in large language models (LLMs) Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023); Feng et al. (2023) have significantly improved the detection of semantically distinct out-of-distribution (far-OOD) intents Liu et al. (2023). However, identifying semantically similar (near-OOD) intents continues to pose significant challenges Fort et al. (2021); Liu et al. (2023). Given its critical importance in practical applications, our study concentrates on near-OOD detection, specifically in a few-shot scenario Jiang et al. (2024); Zhang et al. (2022b). Particularly, in the few-shot learning context, the training model can receive only a few in-distribution (ID) examples per class. Obtaining discriminative information from limited ID examples for OOD detection is inherently challenging Zhang et al. (2020a). Previous strategies have involved using extra unlabeled data Wang et al. (2023) or generating pseudo-ID examples Zhan et al. (2022) to increase training samples. Instead, we focus on leveraging the inner knowledge of LLMs as a supplement without dataset expansion.